# Muchisim: A Simulation Framework for Design Exploration of Multi-Chip Manycore Systems

### v1

@@ III-B Describing and Mapping an Application
The code for an application and its potential DUT compile-time configurations are described as C header files in a structured manner using a template that defines a series of functions. The file for the application to simulate (src/apps) is included via preprocessor macros so that the simulator code is compiled together with the application code.
@@
Configuration functions: The functions with the prefix config_ are invoked only once at the beginning of the simulation. They allow overriding software parameters of the DUT for which the configuration files have set default values. These parameters include the sizes of the queues per task, as well as the parameters for prioritizing the scheduling of certain task IDs. There are, however, other software parameters of the DUT are macros that need to be set when compiling the simulator. The reason why some parameters are macros (instead of being defined in the configuration file) is that they activate different parts of the simulator code that are evaluated every clock cycle. In other words, Muchisim uses pre-processor macros to save branches and improve performance.
@@
Address space and dataset layout: By default, the memory address space is distributed equally and contiguously across the tiles of the DUT. Then, the dataset arrays for a given application are allocated and mapped in memory (at config_app) depending on how a user or mapping technique [95] decides so. On our application suite, the dataset is scattered so that each tile has an equal chunk of it.
@@
Message-triggered tasks (MTTs): The functions associated with an identifier (ID), e.g., task1, describe the different types of MTTs that are invoked upon receiving a message into their corresponding input queue (IQ). Defining these functions enables simulating communication primitives like message passing, active messages, and non-interrupting remote-procedure calls (RPCs). Tasks can invoke other tasks by placing messages into another task’s IQ (if the task is going to be executed on the same PE) or into a channel queue (CQ) if the task is meant to be executed on a different PE. This section later elaborates on the relation between queues and tasks.
@@
Init task: The function with the suffix _init is invoked once at the beginning of each kernel. The kernel_count variable can be used to distinguish behavior between different kernels. Multiple kernels can be simulated in sequence (with global barrier synchronization between them) to compose an application. The combination of init and MTTs per kernel enables simulating a variety of programming models. However, recursion is not supported. On the one extreme, the init task can be empty if the entire kernel only comprises MTTs invoking each other, starting from a seed invocation coming from the DUT’s host. On the other extreme, a kernel can be entirely included in the init task (not invoking remote tasks).
@@
All task functions include the identifier of the PE for which the task is being simulated (and the size of the grid of PEs is macro), so kernels can be parallelized across PEs similarly as with pthreads. section III-C describes how Muchisim iterates over every PE to check whether there are tasks ready to execute.
Association of queues and tasks: Each MTT is associated with an IQ matching its ID. If a task wants to invoke other tasks, it can do so by placing messages into the IQ of the target task ID on the same PE or into the CQs. Each CQ drains into its associated logical channel in the network, and each channel is associated with an IQ ID.
Task dependencies: To avoid network deadlocks, loops in the association between MTTs are not allowed (e.g., task1 invoking task2 and task2 invoking task1). In other words, the dependency chain must end on a leaf task (a task not invoking other tasks); a new phase of computation can start after a local or global barrier. Recall that it is not compulsory for a kernel to use MTTs, but they are the mechanism for the simulator to support message passing and RPCs.
@@ III-C Simulating the Application Runtime
Although the tiles may be organized hierarchically, for parallelization purposes, the simulator considers a global grid of tiles, of which each host thread simulates one or more columns. There are two types of threads, the ones that simulate the TSU and the PU tasks (execution threads) and the ones that simulate the network (router threads).
@@
The execution threads iterate over every tile in their slice of the grid of tiles, checking if there are tasks ready to be executed in the tile’s IQs, in addition to the init task. If there are, the TSU selects the next task type to execute based on the configured policy. Then, the function associated with that task is executed, and the task delay is obtained from the latency-instrumented code. The codes of the application suite included in Muchisim are instrumented considering simple in-order PUs, where the delay is the number of instructions executed (CPI=1); for memory operations, Muchisim offers a special dcache function that returns the latency to fetch a given memory address, depending on whether it hits in the data cache, and the configured memory hierarchy. The function call simulating the task may have pushed messages into other IQs or CQs, with the corresponding timestamps, so they cannot be routed until the router thread is executed for that timestamp. Once the task function returns, the clock for that PU advances as many cycles as the task delay, and the execution thread continues evaluating the next PU. The execution thread keeps iterating over the PUs until there are no more tasks to be done, and the network is empty (i.e., no more messages to be routed). This models a hardware-based termination condition based on idleness, and by default, we set its latency to two times the diameter of the network.
@@
The router threads iterate every cycle over every router in their slide of the grid. For every input port, they check whether there is a message to be routed and a free buffer slot in the destination port (the buffer size is also parameterizable). When two or more input ports want to route to the same output, the priority is determined via a round-robin policy (static priority could also be configured). A message is only routable if the timestamp (set at injection and updated every hop) is less or equal to the current router cycle; the timestamps do not exist in the DUT, but they are used to allow PUs and routers to be simulated in parallel. The synchronization between the execution and router threads is done based on their wall clock time, i.e., a router can never be simulated ahead of its tile’s PUs. In addition, a PU cannot start executing a new task until the router thread has caught up with the clock time of the slowest PU in the tile. Thanks to this type of synchronization, Muchisim can seamlessly support different frequencies for the PUs and the network, with any ratio between them.
Frequency: The configuration file of Muchisim distinguishes between the target frequency established at design for the PU and the routers and their operating frequency. The first is going to affect the area in our current model, while the second affects the power consumption, due to the voltage scaling (described in section III-D). The default target and operating frequency for all PUs, NoC and memory is 1GHz.
@@
SRAM model: The default latency, area and energy parameters of the SRAM memories are based on 7nm technology at 1 Ghz [98] (see table I). Unlike the PU and NoC, SRAM memories have by design a narrower operating voltage and frequency range, so we do not support changing that since we only have a reference for our default values. Muchisim models scaling the size of SRAM memories by increasing the number of banks. Only the active banks consume static energy, and we model the energy of the multiplexor tree that selects the bank to grow by 50% at each doubling step. The access latency is also modeled to increase by 1ns at quadrupling steps beyond 512KiB. All these modeling options and values can be changed in the simulator parameters. On the DUT configurations with multiple PUs per tile, the SRAM is shared among them, but bank conflicts are not modeled yet for this case.
@@
DRAM model: Each compute chiplet may be paired with a memory device of a given number of channels. By default, Muchisim considers an HBM2E device with eight 64GB/s channels. On the 2.5D integration shown in fig. 1, the memory controller sits on one edge of the chiplet, adjacent to the closely integrated DRAM device. The bus connecting the tiles to the memory controller is separate from the task-communication network. Since many tiles share each memory channel, the contention is modeled by considering that the memory channel can only take one request per cycle, and keeping the count of the transactions of each channel. For example, if a request is done at cycle $X$, but the memory channel has received $Y$ transactions (where $Y>X$), then the delay if this request is $Y-X+$ the round-trip to the memory channel. Increasing the capacity of HBM increases the cost linearly (see section III-E) but not the device area since it is a 3D stack. The HBM bandwidth available to the processing units can be defined by changing the number of channels or devices integrated with each compute chiplet, or by changing the number of tiles or PUs per tile of these chiplets.
@@ III-D Energy and Area Model
table I summarizes some of the energy, latency, and area parameters for communication links and memory technology, while the full set of default energy, area and performance parameters of Muchisim can be found in our repo[67]. The simulator periodically logs performance and energy during the simulation, in what we call frames, at a rate that can be configured. This name will become more apparent in section III-F, where we describe our visualization tool, and the key goal of frames is allowing the user to see the progress of the simulation (both at the current frame and aggregated). At the end of the simulation, the simulator also outputs a series of averaged metrics like throughput, average power, and network traffic at the different levels of the hierarchy.
@@
Because for many energy and cost parameters, we do not have a ground truth but rather assumptions or estimates, Muchisim allows post-processing a given simulation to re-calculate the energy and cost with different parameters. In addition to the execution log file, Muchisim creates a separate file with counters collected during simulation, such as messages hops and memory accesses at different levels, and instructions executed for each type. The counters file is then provided as an argument to our separate post-processing executable (which uses the same configuration and parameter files as the simulator, with potentially new values), to re-calculate the energy usage of the execution and the DUT area and cost.
Voltage and area scaling (with frequency) are some of the aspects that can be re-evaluated during post-processing. By default, Muchisim simplistically models the area of the PUs and routers to scale by 50% of the increase in their target frequency. This is a simplistic model that could be refined by synthesizing particular RTL components with different target frequencies. For our voltage scaling model, we fit a ridge regression to the frequency and voltage data from the shmoo plots of chips with 5, 7 and 12nm technology [82, 81, 27]. The current model grows voltage with $0.06+0.13*freq+0.06*node$, which could adjusted by adding data to src/voltage_model.py.
@@ III-E Cost Model
We also added a cost model to the simulator to study the cost-effectiveness of different architecture configurations. The cost model—similarly to the energy model—is decoupled from the runtime simulation process. This is useful to study how price variations can change the cost-effectiveness of different DUT configurations.
Silicon cost: By default, we consider 7nm technology for transistors; we assume that a 300mm wafer with this transistor process costs \$6,047 [37]. We obtain the cost per die by dividing the wafer cost by the number of good dies, which we calculate using 0.2mm scribes, 4mm edge loss, and 0.07 defects per $mm^{2}$. We integrate and validate [34] die yield calculations in our cost model using Murphy’s model. When comparing the cost-effectiveness of the simulated architecture, we do not include the Non-Recurring Engineering (NRE) cost of the compute dies since all the options use the same technology.
Packaging cost: We model the cost of the 65nm silicon interposer connecting a compute die with a DRAM device (including bonding) to be 20% of the price of a compute die [93], and the cost of an organic substrate to be 10% of the price of an equal-sized compute die, and the bonding to add an additional 5% overhead [88, 45].
DRAM cost: While the cost of HBM2 is not disclosed, we made an educated guess using public sources [37, 75]. By default, the cost models assumes 7.5\$/GB, which is more affordable than when HBM was released in 2017. One could expect this price to decrease over time as more vendors fabricate HBM or the process matures [54, 46, 62, 74]. Muchisim’s post-processing tool allows evaluating the performance-per-dollar of a given simulation in the light of different DRAM cost scenarios.
@@ III-F Visualization Tools
Figure 2: Animation of the router activity when running BFS with R22 on a Mesh (top), a Torus (middle), and a Torus with Reduction trees (bottom). The left panels show the routing activity and the right panels show the core (PU) activity. No router activity can mean that the router has no messages to route, or that the NoC is clogged and messages are stuck. The animation is composed of snapshots at a rate of a frame every 40 microseconds (this rate is configurable prior to simulation). Since this rate is the same for these plots, the number of frames (50, 28, and 16, from top to bottom) is proportional to the execution time. The animation can be visualized by opening this PDF with Adobe (for convenience, we have also added these animations as gifs in repository[67].) Visualizing the router and core activity simultaneously helps understand the effect of NoC congestion on core utilization. In addition, plotting the destination port collisions helps understand the router activity. We evaluated the version of BFS with barrier synchronization at the end of each epoch to showcase the effect of the network on the tail execution time (3 major epochs can be seen). A finer time resolution allows observing the evolution of the execution in more detail, but it increases the size of the gif.
@@ III-G Benchmark Suite
Datasets: Muchisim includes ten datasets, including six sizes of the RMAT [48] graphs—standard on the Graph500 list [57]—RMAT-16, 21, 22, 25, 26, and 27, which are named after their scale. For example, RMAT-26 (abbreviated as R26 in section IV-B) contains $2^{26}$, i.e., 67M vertices (V) and 1.3B edges (E), and has a memory footprint of 12GB. We also include the real-world graphs from Wikipedia (V=4.2M, E=101M), LiveJournal (V=5.3M, E=79M), Amazon (V=262K, E=1.2M) and Twitter (V=81K, E=2.4M) [49]. For SPMV and SPMM, the graphs are seen as a square sparse matrix with V rows and columns and E non-zero elements. The graphs (as sparse matrices) are stored in the Compressed Sparse Row (CSR) format without any partitioning, i.e, the dataset contains three input arrays, one for the values of the non-zeros, one for the column indices of those non-zeros, and one for the pointers to the beginning of each row in the previous two arrays. For Histogram, the output array is the count of the column indices of the non-zeros.
@@ IV Results
We include additional experiments and case studies, that we do not have the space for in the current paper, in our repo[67], where we explore the impact of (1) the NoC topology, width and frequency, (2) the reduction-tree size (3) the number of PUs per tile and the PU frequency, (4) integrating DRAM or not, which sets the minimum parallelization required for a given dataset, and (5) the size of the tile input and output queues, among others.

### v2

@@ III-B Describing and Mapping an Application
The code for an application and its potential DUT compile-time configurations are described as C/C++ header files in a structured manner using a template that defines a series of functions. The file for the application to simulate (src/apps) is included via preprocessor macros so that the simulator code is compiled together with the application code.
@@
Configuration functions: The functions with the prefix config_ are invoked only once at the beginning of the simulation. They allow overriding software parameters of the DUT for which the configuration files have set default values. These parameters include the sizes of the queues per task, as well as the parameters for prioritizing the scheduling of certain task IDs. Other software parameters of the DUT are macros that need to be set when compiling the simulator. MuchiSim uses pre-processor macros in some places to save branches and improve performance.
@@
Address space and dataset layout: MuchiSim maps the memory address space contiguously such that the PLM of each tile is assigned a chunk of this contiguous space. The dataset arrays for a given application are allocated and mapped to memory across tiles (at config_app) depending on how a user or mapping technique [96] decides so. On our application suite, the dataset is scattered so that each tile has an equal chunk of each data array.
@@
Message-triggered tasks (MTTs): The functions associated with an identifier (e.g., task1) describe the different types of MTTs that are invoked upon receiving a message into their corresponding input queue. Each MTT is associated with an IQ matching its identifier (ID). Defining MTTs enables simulating communication primitives like message passing, active messages, and non-blocking remote procedure calls (RPCs). Tasks can invoke other tasks by placing messages directly into another task’s input queue (IQ)—if the task is going to be executed on the same PU—or into a channel queue (CQ)—if the task is meant to be executed on a different PU. Each CQ drains into its logical channel in the network, and each channel is associated with an IQ ID. To avoid network deadlocks, loops in the association between MTTs are not allowed (e.g., task1 invoking task2 and task2 invoking task1). In other words, the dependency chain must end on a leaf task (a task not invoking other tasks); a new phase of computation can start after a local or global barrier.
@@
Init task: The function with the suffix _init is invoked once at the beginning of each kernel. The kernel_count variable can be used to distinguish behavior between different kernels. Multiple kernels can be simulated in sequence (with global barrier synchronization between them) to compose an application. The combination of the Init task and MTTs per kernel enables the simulation of a variety of programming models. However, recursion is not supported.
@@
Supported parallelization modes: On the one extreme, the Init task can be empty if the entire kernel only comprises MTTs invoking each other, starting from a seed invocation coming from the DUT’s host. On the other extreme, a kernel can be entirely included in the Init task (not invoking remote tasks). All task functions can access the identifier of the PU for which the task is being simulated and the size of the grid of PUs, and thus, kernels can be parallelized across PUs similarly to using pthreads. section III-C describes how MuchiSim iterates over every PU to check whether there are tasks ready to execute.
@@ III-C Simulating the Application Runtime
Although the tiles of the DUT may be organized hierarchically, for parallelization purposes, the simulator considers a global grid of tiles, of which each host thread simulates a slice of one or more columns. There are two types of threads, the ones that simulate the TSU and the PU tasks (execution threads) and the ones that simulate the network (router threads).
@@
The execution threads iterate over every tile in their slice of the grid, checking if there are tasks ready to be executed in the tile’s IQs, in addition to the Init task. If there are, the TSU selects the next task type to execute based on the configured policy. Then, the function associated with that task is executed, and the task delay is obtained from the latency-instrumented code. The codes included in our application suite are instrumented considering the execution of simple in-order PUs. However, users could change that to reflect other PU models. For memory operations, MuchiSim offers a special dcache function that returns the latency to fetch a given memory address, depending on whether it hits in the data cache, and the configured memory system. The function call simulating the task may have pushed messages into other IQs or CQs, with the corresponding timestamps, so they cannot be routed until the router thread is executed for that timestamp. Once the task function returns, the clock for that PU advances as many cycles as the task delay, and the execution thread continues evaluating the next PU. The execution thread keeps iterating over the PUs until there are no more tasks to be done, and the network is empty (i.e., no more messages to be routed). This models a hardware-based termination condition based on idleness, and by default, we set its latency to two times the diameter of the network.
@@
The router threads iterate every cycle over every router in their slide of the grid. For every input port, they check whether there is a message to be routed and a free buffer slot in the destination port (the buffer size is also parameterizable). When two or more input ports want to route to the same output, the priority is determined via a round-robin policy (static priority could also be configured). A message is only routable if the timestamp (set at injection and updated every hop) is less than or equal to the current router cycle; the timestamps do not exist in the DUT, but they are used to allow PUs and routers to be simulated in parallel. The synchronization between the execution and router threads is done based on their wall clock time, i.e., a router can never be simulated ahead of its tile’s PUs. In addition, a PU cannot start executing a new task until the router thread has caught up with the clock time of the slowest PU in the tile. Thanks to this type of synchronization, MuchiSim can seamlessly support different frequencies for the PUs and the network, with any ratio between them.
Frequency: The DUT configuration file in MuchiSim distinguishes between the peak frequency supported by the design and the operating frequency (at which the DUT is evaluated) for the PUs and the NoCs. Peak design frequency is going to affect the silicon area of the design, while the operating frequency affects the power consumption due to voltage scaling. The default model for this is described in section III-D). Peak and operating frequencies can be changed independently for PUs and NoCs, which is useful to evaluate architectures looking to support applications with different compute and network demands. For example, having a higher peak PU frequency allows raising its operating frequency when needed (e.g., to increase compute capacity) while lowering it to save power for applications with low arithmetic intensity. The configuration file has by default a peak and operating frequency of 1 GHz for all components.
@@
SRAM model: The default latency, area and energy parameters of the SRAM memories are based on 7nm technology at 1 Ghz [99] (see table I). Unlike the PU and NoC, SRAM memories have by design a narrower operating voltage and frequency range, therefore we do not support changing these and we only use the default values for which we have a reference (table I). MuchiSim models scaling the size of SRAM memories by increasing the number of banks. Only the active banks consume static energy, and we model the energy of the multiplexor tree that selects the bank to grow by 50% at each doubling step. The access latency is also modeled to increase by 1 nanosecond at quadrupling steps beyond 512 KiB. All these modeling options and values can be changed in the simulator parameters. On the DUT configurations with multiple PUs per tile, the SRAM is shared among them, but bank conflicts are not modeled yet for this case.
@@
DRAM model: Each compute chiplet may be paired with a memory device of a given number of channels. By default, MuchiSim considers an HBM2E device with eight 64GB/s channels. On the 2.5D integration shown in fig. 1, the memory controller sits on one edge of the chiplet, adjacent to the closely integrated DRAM device. The bus connecting the tiles to the memory controller is separate from the task-communication network. Since each memory channel is shared by many tiles, the contention is modeled by imposing that the memory channel can only take one request per cycle, and keeping the count of the transactions of each channel. For example, if a request is done at cycle $X$, but the memory channel has received $Y$ transactions (where $Y>X$), then the delay if this request is $Y-X+$ the round-trip to the memory channel. Increasing the capacity of HBM increases the cost linearly (see section III-E) but not the device area since it is a 3D stack. The HBM bandwidth available to the processing units can be defined by changing the number of channels or devices integrated with each compute chiplet, or by changing the number of tiles or PUs per tile of these chiplets.
@@ III-D Energy and Area Model
table I summarizes some of the energy, latency, and area parameters for communication links and memory technology, while the full set of default energy, area and performance parameters of MuchiSim can be found in our repo[68]. The simulator periodically logs performance and energy during the simulation, in what we call frames, at a rate that can be configured. This name will become more apparent in section III-F, which describes the visualization tools. Recording frames allows the user to observe the progress of the simulation, by visualizing several metrics (per PU or averaged) for each frame and aggregating until that frame. At the end of the simulation, the simulator also outputs statistics over metrics like throughput, average power, and network traffic at different levels of the hierarchy.
@@
Because there is often no ground truth for energy and cost parameters but rather estimations or assumptions, MuchiSim allows post-processing a given simulation to re-calculate the energy and cost with different model parameters. In addition to the execution log file, MuchiSim creates a separate file with many performance counters collected during the simulation process, such as messages hops and memory accesses at different levels, and instructions executed for each type. The counters file is then provided as an argument to our separate post-processing executable—which uses the same configuration and parameter files as the simulator, with potentially new values—to re-calculate the energy usage, and the DUT area and cost.
The model for how area and voltage grows with peak and operating frequency can be re-evaluated during post-processing. By default, MuchiSim simplistically models the area of the PUs and routers to grow by 50% of the increase in their peak frequency. This model can be refined by synthesizing particular RTL components with different peak frequencies and measuring their area. A simulation can be post-processed with a new area model to re-calculate energy and cost.
For our voltage scaling model, we fit a ridge regression to the frequency and voltage data from the shmoo plots of chips with 5, 7 and 12 nm transistor nodes [83, 82, 28]. The current model grows voltage with $0.06+0.13*freq+0.06*node$, which could be adjusted by adding data to src/voltage_model.py.
@@ III-E Cost Model
We also added a fabrication cost model to the simulator to study the cost-effectiveness of different architecture configurations. Similarly to the energy model, the cost model is decoupled from the runtime simulation process, i.e., cost and energy can be re-calculated post-simulation for different parameters. This is useful to study how price variations can change the cost-effectiveness of different DUT configurations.
Silicon cost: By default, we consider 7 nm technology for transistors; we assume that a 300 mm wafer with this transistor process costs \$6,047 [37]. We obtain the cost per die by dividing the wafer cost by the number of good dies, which we calculate using 0.2 mm scribes, 4 mm edge loss, and 0.07 defects per $mm^{2}$. We integrate and validate [34] die yield calculations in our cost model using Murphy’s model. When comparing the cost-effectiveness of the simulated architecture, we do not include the non-recurring engineering cost of the compute dies since all the options use the same technology.
Packaging cost: We model the cost of the 65 nm silicon interposer connecting a compute die with a DRAM device (including bonding) to be 20% of the price of a compute die [94], and the cost of an organic substrate to be 10% of the price of an equal-sized compute die, and the bonding to add an extra 5% overhead [89, 45].
DRAM cost: While the cost of HBM2 is not disclosed, we made an educated guess using public sources [37, 76]. By default, the cost model assumes \$7.5/GB, which is more affordable than when HBM was first released in 2017. One could expect this price to decrease over time as more vendors fabricate HBM or the process matures [55, 46, 63, 75]. MuchiSim’s post-processing tool allows evaluating the performance-per-dollar of a given simulation in the light of different DRAM cost scenarios.
@@ III-F Visualization Tools
Figure 2: Animation of the router activity when running BFS on RMAT-22 for three different NoCs: 2D mesh (top), 2D torus (middle), and 2D torus with reduction trees (bottom). The left panels show the routing activity and the right panels show the PU (core) activity. No router activity can mean that the router has no messages to route, or that the NoC is clogged and messages are stuck. The animation is composed of snapshots at a rate of a frame every 40 microseconds (this rate is configurable in MuchiSim). Since this rate is the same for these plots, the number of frames (50, 28, and 16, from top to bottom) is proportional to the execution time. The animation can be visualized by opening this PDF with Adobe. Visualizing the router and core activity simultaneously helps understand the effect of NoC congestion on core utilization. In addition, plotting the destination-port collisions helps understand the router activity. We evaluated the version of BFS with barrier synchronization at the end of each epoch to showcase the effect of the network on the tail execution time (3 major epochs can be observed). A finer time resolution allows observing the evolution of the execution in more detail, but it increases the size of the GIF.
@@ III-G Benchmark Suite
Datasets: MuchiSim includes ten datasets, including six sizes of the RMAT [48] graphs—standard on the Graph500 list [58]—RMAT-16, 21, 22, 25, 26, and 27, which are named after their scale. For example, RMAT-26 contains $2^{26}$, i.e., 67M vertices (V) and 1.3B edges (E), and has a memory footprint of 12GB. We also include the real-world graphs from Wikipedia (V=4.2M, E=101M), LiveJournal (V=5.3M, E=79M), Amazon (V=262K, E=1.2M) and Twitter (V=81K, E=2.4M) [49]. For SPMV and SPMM, the graphs are seen as a square sparse matrix with V rows and columns and E non-zero elements. The graphs (as sparse matrices) are stored in the Compressed Sparse Row (CSR) format without any partitioning, i.e, the dataset contains three input arrays, one for the values of the non-zeros, one for the column indices of those non-zeros, and one for the pointers to the beginning of each row in the previous two arrays. For Histogram, the output array is the count of the column indices of the non-zeros.
@@ IV Results
Our repo[68] includes additional experiments and case studies that explore the impact of (1) the NoC topology, width and frequency, (2) the reduction-tree size (3) the number of PUs per tile and the PU frequency, (4) integrating DRAM or not, which sets the minimum parallelization required for a given dataset, and (5) the size of the tile input and output queues, among others.